OpenStudy (anonymous):

Find the sum of the first 12 terms of 4 + 12 + 36 + 108 + …

OpenStudy (aravindg):

What do you notice as relation between two successive terms

OpenStudy (tkhunny):

12/4 = 3 36/12 = 3 108/36 = 3

OpenStudy (anonymous):

So the answer would be 3?

OpenStudy (tkhunny):

Did you add anything? Did you add 12 things?

OpenStudy (anonymous):

Im so confused

OpenStudy (tkhunny):

Read the problem statement. What is it you are supposed to do?

OpenStudy (anonymous):

You are supposed to add

OpenStudy (anonymous):

But what do I add?

OpenStudy (tkhunny):

Add the first 12 terms in that series. 4 + 12 + 36 + 108 + … Add the first 2 terms 4 + 12 = 16 Add the first 3 terms 4 + 12 + 36 = 52 Add the first 4 terms 4 + 12 + 36 + 108 = 160 You have two tasks. What ARE the first 12 terms. and Add them up!

OpenStudy (anonymous):

16+52+160=228

OpenStudy (anonymous):

Is that correct?

OpenStudy (tkhunny):

Why are you adding those sums? Find the first 12 terms and add those.

OpenStudy (anonymous):

How do i find the first 12 terms?

OpenStudy (tkhunny):

That is the right question. Go back and read my very first post. That information should tell you how to do it.

OpenStudy (anonymous):

Ok i got 2934

OpenStudy (tkhunny):

Please don't ever do that - just drop numbers. There is just no way for me to know how you obtained that value. Did you copy it out of a book? Did you pull it out of a hat? How did you get that?

OpenStudy (tkhunny):

Have you, at any time, studied the "Geometric Series"?

OpenStudy (anonymous):

I kept adding dividing by 3 thats how i got it

OpenStudy (tkhunny):

That's no good. We have a Geometric Series First Element is 4 Common Ratio is 3 If this information doesn't mean anything to you, then you are in the wrong class or you have been sleeping or you have read none of your course materials. Help me out, here. Do you have ANY familiarity with Geometric Series?